90 men and 60 women were asked if they had watched the latest ‘Expendables’ movie. Altogether 3/5 of the people said Yes.

3/10 of the women said Yes.

What percentage of the men said No? (Show your working)


1
Expert's answer
2021-09-30T06:56:57-0400

total number of people = 90 + 60 = 150. Number of people who said yes = 3/5 of total = 3/5 X 150 = 90. Number of women who said yes = 3/10 of total women = 3/10 X60 = 18.